# Existence and Uniqueness Results for a Mean-field Game of Optimal Investment

### v2

@@ 3. The optimization problem
We divide our analysis into two subsections, the first devoted to the finite time horizon case, the second one to the infinite time horizon case.
@@ 3.1. The finite time horizon case
Let us consider the finite time horizon case, i.e., fix $T<+\infty$. For each fixed $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$, we consider the problem of maximizing the functional
@@
In the next proposition we are going to show important properties of the functional $J_{T,q}$.
Proposition 3.1.
Fix a random variable $\xi$ satisfying Assumption 2.1, ${\mathbf{u}}\in{\mathcal{U}}_{T}$, and $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Then, the functional $J_{T,q}$ defined in (3.1) is finite and verifies
@@
Proof.
Fix $\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty))$, and ${\mathbf{u}}\in{\mathcal{U}}_{T}$. Since the assumptions of Lemma 2.2 are verified, we can use (2.6) and apply the Fubini-Tonelli theorem (note that the integrand is non-negative) to get
@@
It easily follows that $J_{T,q}$ is finite and that Equation (3.3) holds, which also entails the last statement of the Proposition. ∎
@@
The last statement in Proposition 3.1 implies that the value function $V_{T,q}$ depends on the initial condition $\xi$ only through its average. More precisely, if $\xi,\xi^{\prime}$ both verify Assumption 2.1 and are such that ${\mathbb{E}}[\xi]={\mathbb{E}}[\xi^{\prime}]$, then $V_{T,q}(\xi)=V_{T,q}(\xi^{\prime})$.
Thanks to the Proposition above we can find the optimal control for our optimization problem and this, as a byproduct, allows us to explicitly compute the value function.
Theorem 3.3.
Fix a random variable $\xi$ satisfying Assumption 2.1 and $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Then, $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}\in{\mathcal{U}}_{T}$, where $z^{(T,q)}$ is the function defined in (3.4), is an optimal control for problem (3.2), which is deterministic and independent of $\xi$.
@@
Proof.
Fix $\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty))$. From (3.3) it immediately follows that
@@ 3.2. The infinite time horizon case
We analyze now the infinite time horizon case, i.e., we set $T=+\infty$. For each fixed $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,+\infty);(0,+\infty))$, we consider the problem of maximizing the functional
@@
The next result is analogous to Proposition 3.1. The proof proceeds along the same lines and, thus, we omit it.
Proposition 3.4.
Fix a random variable $\xi$ satisfying Assumption 2.1 and ${\mathbf{u}}\in{\mathcal{U}}_{\infty}$. Consider a function $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))$ and such that the function $z^{(\infty,q)}\colon[0,+\infty)\to(0,+\infty)$ given by
$z^{(\infty,q)}_{s}\coloneqq\int_{s}^{+\infty}{\mathrm{e}}^{-(\rho+\delta)(r-s)}q_{r}^{-\beta}\,{\mathrm{d}}r,\quad s\geq 0,$ (3.10)
is bounded on $[0,+\infty)$.
@@
Also in this case (cf. Remark 3.2) the last statement in Proposition 3.4 entails that the value function $V_{\infty,q}$ depends on the initial condition $\xi$ only through its average.
Thanks to Proposition 3.4, also in the infinite time horizon case we can find the optimal control for problem (3.9) and the explicit form of the corresponding value function, as stated in the next theorem. Its proof is omitted, being similar to that of Theorem 3.3.
Theorem 3.6.
Fix a random variable $\xi$ satisfying Assumption 2.1. Consider a function $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))$ and such that the function $z^{(\infty,q)}$, defined in (3.10) is bounded on $[0,+\infty)$.
@@ 4. Existence and uniqueness of equilibria
As usual in mean-field games theory, looking for an equilibrium boils down to finding a fixed point of a suitable map. According to Definition 2.3, in our case this map is $(q_{s})_{s\in[0,T]}\mapsto\left({\mathbb{E}}[X_{s}^{\xi,\widehat{\mathbf{u}}^{(T,q)}}]\right)_{s\in[0,T]}$, where $\widehat{\mathbf{u}}^{(T,q)}$ is the optimal control for the maximization problem studied in Section 3, whose expression is given either in Theorem 3.3, in the case $T<+\infty$, or in Theorem 3.6, in the case $T=+\infty$.
@@
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 2.3, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty)),$ (4.1)
    then $\widehat{q}$ is a solution to the integral equation
@@
    Vice versa, if there exist a unique solution $\widehat{q}$ to (4.2) satisfying (4.1), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(T,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (4.1), where $z^{(T,q)}$ is the function defined in (3.4).
@@
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 2.3, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))\quad\text{and}\quad z^{(\infty,\widehat{q})}\text{ is bounded on }[0,+\infty)\,,$ (4.3)
    where $z^{(\infty,q)}$ is the function defined in (3.10), then $\widehat{q}$ is a solution to the integral equation
@@
    Vice versa, if there exist a unique solution $\widehat{q}$ to (4.4) satisfying (4.3), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(\infty,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (4.3).
Proof.
@@
(i). Let $(\widehat{\mathbf{u}},\widehat{q})$ be an equilibrium of the mean-field game, with $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\widehat{q}\colon[0,T]\to(0,+\infty)$, such that $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Since point (i) of Definition 2.3 holds, from Theorem 3.3 we have that $\widehat{\mathbf{u}}=z^{(T,\widehat{q})}$, for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. $(\omega,s)\in\Omega\times[0,T]$. Combining the expression of $z^{(T,q)}$, given in (3.4), and (2.6), we get
@@
(ii). Let $\widehat{q}$ be the unique solution to (4.2) such that $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Define $\widehat{\mathbf{u}}\coloneqq z^{(T,\widehat{q})}$ according to (3.4). Then, by Theorem 3.3, we know that $\widehat{\mathbf{u}}$ is optimal for the problem of maximizing $J_{T,\widehat{q}}$, i.e., that point (i) of Definition 2.3 holds. We also have that equation (4.5) above is verified, and hence, since we assumed that $\widehat{q}$ solves (4.2), we get that $\widehat{q}_{s}={\mathbb{E}}[X^{\xi,\widehat{\mathbf{u}}}_{s}]$, for all $s\in[0,T]$, i.e., that point (ii) of Definition 2.3 also holds. This means that $(\widehat{\mathbf{u}},\widehat{q})$ is an equilibrium of the mean-field game.
If $(\widetilde{\mathbf{u}},\widetilde{q})$ is any other equilibrium of the mean-field game, such that $\widetilde{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$, then necessarily $\widetilde{\mathbf{u}}=z^{(T,\widetilde{q})}$, for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. $(\omega,s)\in\Omega\times[0,T]$, by Theorem 3.3. Therefore, we can repeat the same argument of the proof of point (i) to get that $\widetilde{q}$ is a solution to (4.2), and hence that $(\widetilde{\mathbf{u}},\widetilde{q})=(\widehat{\mathbf{u}},\widehat{q})$, since we assumed that (4.2) has a unique solution. ∎
@@
Indeed, this equivalence follows from the next two definitions and from Proposition 4.4 below, whose proof is omitted, being a standard result.
@@
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}z_{s}={\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}z_{u}^{-\beta}\,{\mathrm{d}}u,&s\in[0,T],\\ z_{0}=x\,,\end{dcases}$ (4.7)
@@
The notions of classical and mild solution to (4.7) are analogous to those given in Definitions 4.2 and 4.3. Also in this case, a result similar to Proposition 4.4 shows that these two solution concepts are equivalent.
@@
Note that equations (4.6) and (4.7) do not fall into the standard theory of integro-differential equations, as we have an initial-value problem with the integral being backward in time.
Now, we proceed to study existence and uniqueness of a solution to (4.6) or, equivalently, (4.7). As in the previous section, we are going to divide our analysis into two subsections, the first devoted to the finite time horizon case, the second one to the infinite time horizon case.
@@ 4.1. The finite time horizon case
For each fixed $x>0$, there exists a unique solution to the integro-differential equation (4.7) in the set ${\mathcal{C}}_{x}$ defined in (4.13), which is also twice continuously differentiable. Consequently, the integro-differential equation (4.6) admits a unique solution $y$. Such solution belongs to ${\mathrm{C}}^{2}([0,T];(0,+\infty))$ and satisfies
@@
$f_{s}\leq g_{s}\quad\Longrightarrow\quad\Gamma(f)_{s}\geq\Gamma(g)_{s},\quad s\in[0,T],$
@@
$\sup_{f\in{\mathcal{C}}}\sup_{s\in[0,T]}\lvert\Gamma(f)_{s}^{\prime}\rvert=\sup_{f\in{\mathcal{C}}}\sup_{s\in[0,T]}\left\{{\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u\right\}\leq\sup_{s\in[0,T]}\frac{{\mathrm{d}}}{{\mathrm{d}}s}\overline{\alpha}^{x}_{s}<+\infty.$
@@
Case 2. Consider $w_{0}=\widetilde{w}_{0}$ and note that both $w$ and $\widetilde{w}$ are bounded below by the constant $x$, since $z,\widetilde{z}\in{\mathcal{C}}_{x}$. Observe, also, that the function $w\mapsto w^{-\beta}$ is Lipschitz in $[x,+\infty)$, with Lipschitz constant $L\coloneqq x^{-\beta-1}$. Therefore, integrating (4.17) on $[0,s]$ we get
@@
Therefore, equation (4.7) admits a unique solution in ${\mathcal{C}}_{x}$. Finally, from Proposition 4.6, we immediately deduce that also (4.6) has a unique solution in ${\mathrm{C}}^{2}([0,T];(0+\infty))$ satisfying the bounds in (4.14). ∎
@@
Proposition 4.10.
For each fixed random variable $\xi$ satisfying Assumption 2.1, there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ for the mean-field game with finite time horizon, among all equilibria verifying the integrability condition (4.1).
@@
Therefore, $\widehat{q}$ satisfies the integrability condition (4.1) and applying Theorem 4.1-(ii) we get the result. ∎
@@ 4.2. The infinite time horizon case
Let $T=+\infty$ and assume that $\beta<1+\frac{\rho}{\delta}$. Then, for each fixed $x>0$, there exists a unique solution to the integro-differential equation (4.7) in the set ${\mathcal{C}}_{x,\infty}$ defined in (4.22), which is also twice continuously differentiable. Moreover, the integro-differential equation (4.6) admits a unique solution $y$. Such solution belongs to ${\mathrm{C}}^{2}([0,+\infty);(0,+\infty))$ and satisfies
@@
Proof.
We divide the proof as follows. First, we show existence and uniqueness of a solution to (4.7) in the set ${\mathcal{C}}_{x,\infty}$, for each fixed $x>0$. As in the finite time horizon case, the fact that such solution is twice continuously differentiable follows by differentiation of (4.8). Then, it immediately follows from Proposition 4.6 that (4.6) admits a unique solution $y\in{\mathrm{C}}^{2}([0,+\infty);(0,+\infty))$ satisfying the bounds given in (4.23). Finally, we will show that the constant function equal to $y_{\infty}$ is the unique constant solution to (4.6), and that the solutions starting at $x\neq y_{\infty}$ are monotone in time and converge to $y_{\infty}$.
@@
$\leq\dfrac{K({\mathrm{e}}^{\delta(1+\beta)h}-1)}{\delta(1+\beta)(\rho+\delta-\delta\beta)}+\dfrac{K}{\delta(1+\beta)(\rho+2\delta)}[{\mathrm{e}}^{(\rho+2\delta)h}-1],\quad\forall n\geq 1\,.$
Therefore, we obtain that
@@
with $g\in{\mathrm{L}}^{1}_{\rho+2\delta}$. Therefore, the sequence of derivatives $\{\left(z^{(n)}\right)^{\prime}\}$ converges in ${\mathrm{L}}^{1}_{\rho+2\delta}$ to $w^{(\infty)}$. We deduce that the subsequence $\{z^{(n)}\}$ converges in the weighted Sobolev space $W^{1,1}_{\rho+2\delta}$ (cf. [4, Chapter 8, Remark 4]). By completeness of this space, we get that $w^{(\infty)}_{s}=(z^{(\infty)}_{s})^{\prime}$, for a.e. $s\geq 0$. Moreover, using the same argument as in the proof of [4, Theorem 8.2], we get that there exists a function $\overline{z}^{(\infty)}\in{\mathrm{C}}([0,+\infty);(0,+\infty))$ such that $\overline{z}^{(\infty)}_{s}=z^{(\infty)}_{s}$, for almost every $s\geq 0$, and
@@
Note that we only need to consider $\zeta>0$, since solutions to (4.7) are such that their first derivative is strictly positive for all $s\geq 0$. For each fixed $x>0$ and $\zeta>0$, (4.26) has a unique solution $z^{x,\zeta}$ on the maximal interval $[0,\tau^{*}(x,\zeta))$, with $\tau^{*}(x,\zeta)\leq+\infty$.
@@
Let us fix $x>0$. Standard results (see, e.g., [11, Theorem 7,5, Chapter 1]) ensure that the solution $z^{x,\zeta}$ to (4.26) and its first- and second-order derivatives with respect to the time variable are differentiable with respect to $\zeta$ and that the order of differentiation can be exchanged.
@@
$\begin{dcases}\widehat{z}^{\prime\prime}_{s}(\zeta)=(\rho+2\delta)\widehat{z}^{\prime}_{s}(\zeta)+\beta{\mathrm{e}}^{\delta(1+\beta)s}z_{s}(\zeta)^{-\beta-1}\widehat{z}_{s}(\zeta),&s\geq 0,\\ \widehat{z}_{0}(\zeta)=0\,,\\ \widehat{z}^{\prime}_{0}(\zeta)=1\,,\end{dcases}$ (4.27)
where $z_{s}(\zeta)\coloneqq z_{s}^{x,\zeta}$. Let
@@
$\begin{dcases}\widehat{z}^{\prime\prime}_{s}(\zeta)\geq(\rho+2\delta)\widehat{z}^{\prime}_{s}(\zeta),&s\geq 0,\\ \widehat{z}^{\prime}_{0}(\zeta)=1\,,\end{dcases}$ (4.28)
@@
$z_{s}(b)\geq z_{s}(\overline{\zeta})+(b-\overline{\zeta})[{\mathrm{e}}^{(\rho+2\delta)s}-1]\geq x+(b-\overline{\zeta})\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
@@
$z_{s}(a)\leq z_{s}(\overline{\zeta})-(\overline{\zeta}-a)[{\mathrm{e}}^{(\rho+2\delta)s}-1]\leq\overline{\alpha}_{s}^{x,\infty}-(\overline{\zeta}-a)\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
@@
The second case leads to a contradiction as well. Indeed, we would have two different solutions in the interval $[0,s_{0}]$ to the Cauchy problem (4.29), with terminal conditions $y_{s_{0}}=y_{\infty}$ and $y^{\prime}_{s_{0}}=0$, which clearly admits unique solution, that is, $y_{s}=y_{\infty}$, for all $s\geq 0$.
@@
Proposition 4.13.
For each fixed random variable $\xi$ satisfying Assumption 2.1, there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ for the mean-field game with infinite time horizon, among all equilibria verifying condition (4.3).
@@
Proof.
Let $\widehat{q}$ be the unique solution to (4.6), with $x={\mathbb{E}}[\xi]$. We need to check that condition (4.3) is satisfied. We have the following three cases.

### v3

@@ 3. The optimization problem
We divide our analysis into two subsections, the first devoted to the finite time horizon case, the second one to the infinite time horizon case. Since an additional technical condition is required to solve the infinite time horizon case, we discuss the two problems separately, for the sake of clarity.
@@ 3.1. The finite time horizon case
Let us consider the finite time horizon case, i.e., fix $T<+\infty$. We introduce the set
${\mathcal{Q}}_{T}\coloneqq\{q\colon[0,T]\to(0,+\infty)\text{ s.t. }q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))\}\,.$ (3.1)
For each fixed $q\in{\mathcal{Q}}_{T}$, we consider the problem of maximizing the functional
@@
In the next Lemma we are going to show important properties of the functional $J_{T,q}$.
Lemma 3.1.
Fix a random variable $\xi$ satisfying Assumption 2.1, ${\mathbf{u}}\in{\mathcal{U}}_{T}$, and $q\in{\mathcal{Q}}_{T}$. Then, the functional $J_{T,q}$ defined in (3.2) is finite and verifies
@@
Dimostrazione.
Fix $\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty))$, ${\mathbf{u}}\in{\mathcal{U}}_{T}$, and $q\in{\mathcal{Q}}_{T}$. Since the assumptions of Lemma 2.2 are verified, we can use (2.4) and apply the Fubini-Tonelli theorem (note that the integrand is non-negative) to get
@@
It easily follows that $J_{T,q}$ is finite and that Equation (3.4) holds, which also entails the last statement of the Lemma. ∎
@@
The last statement in Lemma 3.1 implies that the value function $V_{T,q}$ depends on the initial condition $\xi$ only through its average. More precisely, if $\xi,\xi^{\prime}$ both verify Assumption 2.1 and are such that ${\mathbb{E}}[\xi]={\mathbb{E}}[\xi^{\prime}]$, then $V_{T,q}(\xi)=V_{T,q}(\xi^{\prime})$.
Thanks to the Lemma above we can find the optimal control for our optimization problem and this, as a byproduct, allows us to explicitly compute the value function.
Proposition 3.3.
Fix a random variable $\xi$ satisfying Assumption 2.1 and $q\in{\mathcal{Q}}_{T}$. Then, $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}\in{\mathcal{U}}_{T}$, where $z^{(T,q)}$ is the function defined in (3.5), is an optimal control for problem (3.3), which is deterministic and independent of $\xi$.
@@
Dimostrazione.
Fix $\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty))$ $q\in{\mathcal{Q}}_{T}$. From (3.4) it immediately follows that
@@ 3.2. The infinite time horizon case
We analyze now the infinite time horizon case, i.e., we set $T=+\infty$. Let us define, for any given measurable $q\colon[0,+\infty)\to(0,+\infty)$ the function
$z^{(\infty,q)}_{s}\coloneqq\int_{s}^{+\infty}{\mathrm{e}}^{-(\rho+\delta)(r-s)}q_{r}^{-\beta}\,{\mathrm{d}}r\,,\quad s\geq 0\,.$ (3.9)
We introduce the set
${\mathcal{Q}}_{\infty}\coloneqq\{q\colon[0,+\infty)\to(0,+\infty)\text{ s.t. }q^{-\beta}\in{\mathrm{L}}^{1}((0,+\infty);(0,+\infty))\text{ and }z^{(\infty,q)}\text{ is bounded}\}\,.$ (3.10)
Note that in the definition of $\mathcal{Q}$$_{\infty}$ we require an additional technical condition with respect to the definition of ${\mathcal{Q}}_{T}$ in the finite time horizon case. This is needed to prove the results of this subsection, namely, Lemma 3.4 and Proposition 3.6.
For each fixed $q\in{\mathcal{Q}}_{\infty}$ we consider the problem of maximizing the functional
@@
The next result is analogous to Lemma 3.1. The proof proceeds along the same lines and, thus, we omit it.
Lemma 3.4.
Fix a random variable $\xi$ satisfying Assumption 2.1, ${\mathbf{u}}\in{\mathcal{U}}_{\infty}$, and $q\in{\mathcal{Q}}_{\infty}$.
@@
Also in this case (cf. Remark 3.2) the last statement in Lemma 3.4 entails that the value function $V_{\infty,q}$ depends on the initial condition $\xi$ only through its average.
Thanks to Lemma 3.4, also in the infinite time horizon case we can find the optimal control for problem (3.12) and the explicit form of the corresponding value function, as stated in the next Proposition. Its proof is omitted, being similar to that of Proposition 3.3.
Proposition 3.6.
Fix a random variable $\xi$ satisfying Assumption 2.1 and $q\in{\mathcal{Q}}_{\infty}$.
@@ 4. Existence and uniqueness of equilibria
As usual in mean-field games theory, looking for an equilibrium boils down to finding a fixed point of a suitable map. According to Definition 2.3, in our case this map is $(q_{s})_{s\in[0,T]}\mapsto\left({\mathbb{E}}[X_{s}^{\xi,\widehat{\mathbf{u}}^{(T,q)}}]\right)_{s\in[0,T]}$, where $\widehat{\mathbf{u}}^{(T,q)}$ is the optimal control for the maximization problem studied in Section 3, whose expression is given either in Proposition 3.3, in the case $T<+\infty$, or in Proposition 3.6, in the case $T=+\infty$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Dimostrazione.
@@
[…]
@@
[…]
@@
Indeed, this equivalence follows from the next definition and from Lemma 4.3 below, whose proof is omitted, being a standard result, see, e.g., [26, Chapter 1].
@@
[…]
@@
The notions of classical and mild solution to (4.5) are analogous to those given in Definition 4.2. Also in this case, a result similar to Lemma 4.3 shows that these two solution concepts are equivalent.
@@
Note that equations (4.4) and (4.5) do not fall into the standard theory of integro-differential equations (see, e.g. [23, 26]), as we have an initial-value problem with the integral being backward in time.
Now, we proceed to study existence and uniqueness of a solution to (4.4) or, equivalently, (4.5). As in the previous section, we are going to divide our analysis into two subsections, the first devoted to the finite time horizon case, the second one to the infinite time horizon case. In particular, to prove the well-posedness of the integro-differential equation (4.4) in the finite time horizon case, we are going to rely on a fixed point argument to show existence of a solution, and on an ad-hoc argument based on the structure of the equation to show its uniqueness. This result will be used in the infinite time horizon case to show existence of a solution to (4.4), by constructing a candidate one as a sequence of solutions to the integro-differential equation with finite time horizon. Then, we prove uniqueness of the solution and its convergence to a steady state using arguments that rely on the analysis of its properties.
@@ 4.1. The finite time horizon case
For each fixed $x>0$, there exists a unique solution to the integro-differential equation (4.5) in the set ${\mathcal{C}}_{x}$ defined in (4.11), which is also twice continuously differentiable. Consequently, the integro-differential equation (4.4) admits a unique solution $y\in{\mathrm{C}}^{2}([0,T];(0,+\infty))$ satisfying the bounds
@@
$f_{s}\geq g_{s}\quad\Longrightarrow\quad\Gamma(f)_{s}\leq\Gamma(g)_{s},\quad s\in[0,T],$
@@
$\sup_{f\in{\mathcal{C}}}\sup_{s\in[0,T]}\lvert\Gamma(f)_{s}^{\prime}\rvert=\sup_{f\in{\mathcal{C}}}\sup_{s\in[0,T]}\left\{{\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u\right\}\leq\sup_{s\in[0,T]}{\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}\overline{\alpha}^{x}_{s}\right|}<+\infty.$
In fact, since $f_{s}^{-\beta}\leq(\underline{\alpha}^{x}_{s})^{-\beta}=x^{-\beta}$, for all $s\in[0,T]$, we have that, if $\beta\neq 1+\frac{\rho}{\delta}$,
$0$ $\leq{\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u$
$\leq\dfrac{x^{-\beta}}{\rho+\delta-\delta\beta}\left\{{\mathrm{e}}^{\delta(1+\beta)s}-{\mathrm{e}}^{-(\rho+\delta-\delta\beta)T}{\mathrm{e}}^{(\rho+2\delta)s}\right\}=\frac{{\mathrm{d}}}{{\mathrm{d}}s}\overline{\alpha}^{x}_{s},\quad s\in[0,T],$
while if $\beta=1+\frac{\rho}{\delta}$,
$0\leq{\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u\leq x^{-\beta}(T-s){\mathrm{e}}^{(\rho+2\delta)s}=\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}\overline{\alpha}^{x}_{s}\right|,\quad s\in[0,T].$
@@
Case 2. Consider $w_{0}=\widetilde{w}_{0}$ and note that both $w$ and $\widetilde{w}$ are bounded below by the constant $x$, since $z,\widetilde{z}\in{\mathcal{C}}_{x}$. Observe, also, that the function $w\mapsto w^{-\beta}$ is Lipschitz in $x,+\infty)$, with Lipschitz constant $L\coloneqq x^{-\beta-1}$. Therefore, integrating (4.15) on $[0,s]$ we get
@@
Therefore, equation (4.5) admits a unique solution in ${\mathcal{C}}_{x}$. Finally, from Lemma 4.5, we immediately deduce that also (4.4) has a unique solution in ${\mathrm{C}}^{2}([0,T];(0+\infty))$ satisfying the bounds in (4.12). ∎
@@
Theorem 4.9.
[…]
@@
Therefore, $\widehat{q}\in{\mathcal{Q}}_{T}$ and applying Theorem 4.1-(ii) we get the result. ∎
@@ 4.2. The infinite time horizon case
As the notation suggests, the function $\overline{\alpha}^{x,\infty}$ given in (4.19) is the pointwise limit, as $T\to+\infty$, of the sequence $\{\overline{\alpha}^{x,T}\}_{T>0}$, where for each $T>0$ the function $\overline{\alpha}^{x,T}$ has the expression given in (4.9).
@@
Let $T=+\infty$ and assume that $\beta<1+\frac{\rho}{\delta}$. Then, for each fixed $x>0$, there exists a unique solution to the integro-differential equation (4.5) in the set ${\mathcal{C}}_{x,\infty}$ defined in (4.20), which is also twice continuously differentiable. Moreover, the integro-differential equation (4.4) admits a unique solution $y\in{\mathrm{C}}^{2}([0,+\infty);(0,+\infty))$ satisfying the bounds
@@
Dimostrazione.
We divide the proof as follows. First, we show existence and uniqueness of a solution to (4.5) in the set ${\mathcal{C}}_{x,\infty}$, for each fixed $x>0$. As in the finite time horizon case, the fact that such solution is twice continuously differentiable follows by differentiation of (4.6). Then, it immediately follows from Lemma 4.5 that (4.4) admits a unique solution $y\in{\mathrm{C}}^{2}([0,+\infty);(0,+\infty))$ satisfying the bounds given in (4.21). Finally, we will show that the constant function equal to $y_{\infty}$ is the unique constant solution to (4.4), and that the solutions starting at $x\neq y_{\infty}$ are monotone in time and converge to $y_{\infty}$.
@@
$\leq\dfrac{K({\mathrm{e}}^{\delta(1+\beta)h}-1)}{\delta(1+\beta)(\rho+\delta-\delta\beta)}+\dfrac{K}{\delta(1+\beta)(\rho+2\delta)}[{\mathrm{e}}^{(\rho+2\delta)h}-1],\quad\forall n\geq 1\,$,
where in the last inequality we used the condition $\beta<1+\frac{\rho}{\delta}$. Therefore, we obtain that
@@
[…]
@@
Note that we only need to consider $\zeta>0$, since solutions to (4.5) are such that their first derivative is strictly positive for all $s\geq 0$. For each fixed $x>0$ and $\zeta>0$, (4.24) has a unique solution $z^{x,\zeta}$ on the maximal interval $0,\tau^{*}(x,\zeta))$, with $\tau^{*}(x,\zeta)\leq+\infty$.
@@
Let us fix $x>0$. Standard results (see, e.g., [14, Theorem 7,5, Chapter 1]) ensure that the solution $z^{x,\zeta}$ to (4.24) and its first- and second-order derivatives with respect to the time variable are differentiable with respect to $\zeta$ and that the order of differentiation can be exchanged.
@@
[…]
where $z_{s}(\zeta)\coloneqq z_{s}^{x,\zeta}$.
Let
@@
[…]
@@
$z_{s}(b)\geq z_{s}(\overline{\zeta})+(b-\overline{\zeta}){\frac{1}{\rho+2\delta}}[{\mathrm{e}}^{(\rho+2\delta)s}-1]\geq x+(b-\overline{\zeta})\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
@@
$z_{s}(a)\leq z_{s}(\overline{\zeta})-(\overline{\zeta}-a){\frac{1}{\rho+2\delta}}[{\mathrm{e}}^{(\rho+2\delta)s}-1]\leq\overline{\alpha}_{s}^{x,\infty}-(\overline{\zeta}-a)\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
@@
The second case leads to a contradiction as well. Indeed, we would have two different solutions in the interval $0,s_{0}]$ to the Cauchy problem (4.27), with terminal conditions $y_{s_{0}}=y_{\infty}$ and $y^{\prime}_{s_{0}}=0$, which clearly admits unique solution, that is, $y_{s}=y_{\infty}$, for all $s\geq 0$.
@@
Theorem 4.12.
[…]
@@
Dimostrazione.
Let $\widehat{q}$ be the unique solution to (4.4), with $x={\mathbb{E}}[\xi]$. We need to check that $\widehat{q}\in{\mathcal{Q}}_{\infty}$. We have the following three cases.